# Never-ending Search for Innovation

### v1

@@ 1 Introduction
The above example has two important features. First, given a set of potential vectors, discovering that one is not responsible for transmission is informative about the probability that a remaining vector is how the disease actually spreads. For instance, if the parasite is not found in the soil, the probability that it will be found in water increases. Second, the cost of testing different vectors varies greatly. For example, examining soil samples is cheap, while dissecting living hosts is costly.
@@
We propose a model exhibiting both of the above features, informational externalities and heterogeneous projects. In doing so, we go beyond the existing literature and we find that, if the search starts, it does not end until the answer is found. In particular, the search will go on forever with strictly positive probability. This is in sharp contrast with the existing results in the literature, where the search tends to end in finite time even if the optimal search rule prescribes to continue searching forever. This can be seen in McCardle et al. (2018), who study monopolistic search in an infinite-horizon, continuous-time bandit model. In their model, depending on the arrival rate of successes and failures, the optimal stopping time is finite or infinite. However, even when the optimal stopping time is infinite, the search will end in finite time with a probability going to 1, as eventually either a success or a failure will realize. [^2] A similar observation holds true in the seminal search models (without informational externalities) such as McCall (1965) and Morgan (1983). See Lippman and McCall (1976) and Chade et al. (2017) for surveys, or for a more recent example see Benkert et al. (2018). In this class of models, even if the optimal search rule prescribes searching forever, the probability that search continues forever will converge to zero, thus ending search in finite time.
@@
Instead of the bandit framework, we study a model where the agent is faced with a set of research projects of which only one has the potential to lead to the breakthrough. Like the bandit framework, our model also features learning, as failure from one set of projects is informative about the chances that a success will be found in the remaining set. Our model is closely related to de Roos et al. (2018), who also consider a model where an agent searches in a set of research projects. Importantly, they assume that all projects are homogeneous, and that a successful project exists with certainty. When this is the case, then the optimal search prescribes that the entire set of available projects is examined in finite time, causing the search to also end in finite time. [^3] Using a similar model as de Roos et al. (2018), Matros and Smirnov (2011) and Matros and Smirnov (2016) focus on the strategic interaction between multiple searchers. In extensions, they consider heterogeneous projects (in terms of success probabilities rather than costs) and allow for no successful project to exist. Nevertheless, they always conclude that search ends in finite time, contrasting with our findings. We instead assume that projects are sufficiently heterogeneous, so that a rational agent optimally never examines the entire set of available projects. Paired with the possibility that no project could be successful, we obtain the contrasting result that optimal search may never end.
@@ 4 Conclusion
This is not merely a technical point, but rather has an important policy implication for the question of when to optimally abandon a (research) project. McCardle et al. (2018) note that, in practice, search often lasts longer than what is perceived as optimal and suggest that one should set an “alarm clock” to ensure a timely end to the search effort. Our results suggest that this might not always be the appropriate policy, and that a research project may optimally need to continue for longer than intuition alone would recommend. According to our results, an outside decision maker (who does not fully account for the informational content of the failures that the agent has experienced) is especially likely to perceive the search as going beyond the optimal threshold. This is particularly relevant in environments with informational spillovers (so that failure today is informative of success chances tomorrow) and adjustable research intensity, which we would argue are more common than not in research. Empirically, our result also highlight the value of collecting data on research intensity, in addition to data on whether research is active or not.

### v2

@@ 1 Introduction
The above example has two important features. First, given a set of potential vectors, discovering that one is not responsible for transmission is informative about the probability that a remaining vector is how the disease actually spreads. For instance, if the parasite is not found in the soil, the researcher may update her belief that the available methods are unable to detect the parasite, leading her to become more pessimistic about the probability of finding the parasite in, for example, water. At the same time, conditional on being able to detect the parasite and the parasite being present in one environment only, then failing to detect the parasite in soil increases the belief of the researcher that she will detect it in water. Second, the cost of testing different vectors varies greatly. For example, examining soil samples is cheap, while dissecting living hosts is costly.
@@
We propose a model exhibiting both of the above features, informational externalities and heterogeneous projects. In doing so, we go beyond the existing literature and we find that, if the search starts, it does not end until the answer is found. In particular, the search will go on forever with strictly positive probability. This is in sharp contrast with the existing results in the literature, where the search tends to end in finite time even if the optimal search rule prescribes to continue searching forever. [^2] This holds true in the seminal search models (without informational externalities) such as McCall (1965) and Morgan (1983). See Lippman and McCall (1976) and Chade et al. (2017) for surveys, or for more recent examples see Benkert et al. (2018) and McCardle et al. (2018). In this class of models, even if the optimal search rule prescribes searching forever, the probability that search continues forever will converge to zero, thus ending search in finite time.
@@
We study a model where the agent is faced with a set of research projects of which only one has the potential to lead to the breakthrough. Importantly, our model features learning, as failure from one set of projects is informative about the chances that a success will be found in the remaining set. Our model is closely related to de Roos et al. (2018), who also consider a model where an agent searches in a set of research projects. Importantly, they assume that all projects are homogeneous, and that a successful project exists with certainty. When this is the case, then the optimal search prescribes that the entire set of available projects is examined in finite time, causing the search to also end in finite time. [^3] Using a similar model as de Roos et al. (2018), Matros and Smirnov (2011) and Matros and Smirnov (2016) focus on the strategic interaction between multiple searchers. In extensions, they consider heterogeneous projects (in terms of success probabilities rather than costs) and allow for no successful project to exist. Nevertheless, they always conclude that search ends in finite time, contrasting with our findings. We instead assume that projects are sufficiently heterogeneous, so that a rational agent optimally never examines the entire set of available projects. Paired with the possibility that no project could be successful, we obtain the contrasting result that optimal search may never end.
@@
It is instructive to compare our framework with the class of bandit models, which are often used to model dynamic innovation. In a standard bandit model, [^4] As an example of such a model used in the context of innovation see Bergemann and Hege (2005). the searcher faces a project (or an arm) with an unknown return. The project is either good, so that investing in it yields a success with some probability, or bad, so that any investment is wasted. In these models, the absence of a success is informative and leads to increasing pessimism: as failures accumulate, the searcher lowers her belief about the probability that she is facing a good project, until she becomes so pessimistic that she stops searching, which happens in finite time. Our model features (infinitely) many heterogeneous projects, and the structure of the model is such that becoming more pessimistic about one project necessarily leads the searcher to become more optimistic about the unexplored projects. This positive informational spillover across projects, together with project heterogeneity, generates a never-ending search in our model, while a standard bandit model results in search ending in finite time.
@@ 4 Conclusion
This is not merely a technical point, but rather has an important policy implication for the question of when to optimally abandon a (research) project. McCardle et al. (2018) note that, in practice, search often lasts longer than what is perceived as optimal and suggest that one should set an “alarm clock” to ensure a timely end to the search effort. However, our results indicate that such a policy may be short-sighted. What appears to be excessive search from an external perspective may in fact be optimal, as the agent internalizes the informational value of continued failures and adjusts search intensity accordingly. An outside decision-maker, who does not fully account for these learning dynamics, may wrongly conclude that search has gone on “too long” and intervene prematurely. This is particularly relevant in environments with informational spillovers (so that failure today is informative of success chances tomorrow) and adjustable research intensity, which we would argue are more common than not in research. Moreover, our result suggests that the debate over whether research lasts “too long” or stops “too soon” is too simplistic. Rather than a binary choice, search intensity can decline over time while still continuing indefinitely. To address this empirically, our finding highlights the need to track not just whether research is ongoing, but how its intensity evolves.